# Iterative Deepening Sampling for Large Language Models

### v1

@@ 4 Iterative Deepening Sampling
Before going into the details, a key challenge in ID-sampling is that budget control occurs before each manually triggered self-evaluation and self-correction step without explicitly analyzing the actual generated responses. This can lead to unnecessary iterations, potentially increasing computational costs. However, due to the design of our ID-sampling method, we establish important theoretical guarantees. In particular, we provide a bound on the total number of tokens generated before reaching the final answer, as formalized in the following theorem.
@@
Suppose the final answer obtained through ID-Sampling needs a budget of $L$. Then the total number of budget used is no more than $\frac{\gamma*L}{\gamma+1}$.
@@
Note that the budget for each iteration of generation is a geometric sequence with a common ratio of $\gamma$, the theorem is a direct conclusion by using a summarization of the geometric sequence.
@@
[…]
Leveraging modern serving engines to efficiently allocate the KV-cache and execute inference in batches using data parallelism and pipeline parallelism significantly enhances the efficiency of Iterative Deepening (ID) sampling. By setting a maximum generation token limit in the sampling parameters, the computational cost scales linearly with generation time, ensuring that ID-sampling remains efficient even as the number of iterations increases.
@@ 5.1 Experiment Setup
In our experiments, we evaluate Iterative Deepening (ID) sampling on a set of mathematical problems. Given the strong problem-solving capabilities of modern models, which inherently possess self-evaluation and self-correction mechanisms, we focus on more challenging benchmarks and omit relatively easier datasets such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). Instead, we directly assess performance on competition-level and Olympiad-level benchmarks, including MATH-500 (Lightman et al., 2023) and AIME-24. Specifically, MATH-500 is a curated subset of 500 problems from the original MATH dataset, designed to evaluate process reward models trained for mathematical reasoning. Meanwhile, AIME (American Invitational Mathematics Examination) is a highly competitive exam aimed at identifying top-performing high school students in the U.S. The AIME-24 dataset comprises 30 problems sourced from the 2024 AIME I and II exams, providing a rigorous benchmark for assessing reasoning capabilities under competition-level constraints. For both datasets, answers are evaluated using symbolic-based checkers to ensure that all mathematically equivalent responses are recognized as correct.
For our experiments, we primarily use LLAMA-3.1-1B-Instruct (Dubey et al., 2024) and Phi-4 (Abdin et al., 2024) before transitioning to the latest DeepSeek-R1-Distill-Qwen-7B (DeepSeek-AI et al., 2025). This allows us to investigate how the strong built-in self-correction capabilities of recently released R1 models influence the performance of our algorithm. For reward models, we employ the newly released Qwen-2.5-Math-PRM-7B (Zhang et al., 2025), which demonstrates superior accuracy compared to other available models. This choice ensures a robust evaluation of reasoning performance in our study.
For baselines, we compare our approach against vanilla Best-of-N (BoN) sampling. For evaluation, we report three key metrics:
@@
For both BoN and majority voting, a single aggregated answer is compared against the ground-truth solution to measure accuracy. Note that when $N=1$, BoN is necessarily reporting the Pass@1 score.
All experiments are conducted on an NVIDIA A100 GPU (80GB memory) using the vLLM serving engine (Kwon et al., 2023). By default, we use a carefully selected set of hyperparameters to balance efficiency and performance:
  • Based on the typical sampling budget used in training (10–40), we set the maximum number of N to 32 for Best-of-N (BoN) sampling.
  • For ID-sampling with BoN, we set the initial budget $B_{0}$ as 256 and terminate when the maximum sequence length reaches 8192 for Deepseek-R1-distill-Qwen-7B and 4096 otherwise.
  • By default, we use $\gamma=2.0$ for ID-sampling, and we will later provide an ablation study results on this.
@@ 5.2 Results
Before evaluating the pass rate, we first analyze the runtime overhead of ID-sampling. We observe that ID-sampling requires approximately 1.6–1.9× the runtime of a baseline approach without the trigger sentence. Given that this additional computational cost could instead be allocated to generating more responses and selecting the best one, we present our results in terms of an equivalent N. Specifically, if the original results correspond to Bo8, we report them as equivalent $N=16$, as ID-sampling consistently completes within twice the runtime of the original method.
@@ 5.2.1 Math-500
Figure 2: Results for Phi4 on Math-500 dataset. The x-axis is the equivalent N after considering the additional time used by ID-sampling.
Figure 3: Results for Deepseek-R1-Distill-Qwen-7B on Math-500 dataset. The x-axis is the equivalent N after considering the additional time used by ID-sampling.
Figure 4: Results for Llama-3.1-1B-Instruct on Math-500. The x-axis is the equivalent N after considering the additional time used by ID-sampling.
The MATH-500 dataset is a curated subset of the MATH dataset, selected to retain its more challenging problems. Originally used by OpenAI to train process reward models (PRMs) (Lightman et al., 2023), this dataset benefits from the high accuracy of recently released reward models. As a result, in this setting, Best-of-N (BoN) sampling can serve as a reliable sampling method even in the absence of a ground-truth checker.
We present our results for different models on Math-500 in Fig.2, 3, 4. Across all scenarios, we observe that ID-sampling provides a significant improvement at the same $N$ and remains effective in most scenarios even when considering the additional computational cost as the equivalent $N$. However, the majority voting score from ID-sampling does not always outperform. The performance gap in majority voting can be attributed to the relative self-correction capabilities of different models and their built-in ability to trigger self-correction. For models with limited self-correction capabilities, manually inserting a trigger sentence, as done in ID-sampling, can sometimes degrade performance, potentially turning previously correct responses into incorrect ones. Because of this, in our setting where N responses are sampled independently, ID-sampling results in a more uniform response distribution. As a consequence, incorrect answers now have a higher probability of being mistakenly selected as the final output under majority voting. While BoN mitigates this issue by leveraging a separate reward model, majority voting relies solely on frequency-based aggregation, making it less effective in such cases.
Our results also indicate that the performance gap between ID-sampling and vanilla sampling increases as model capabilities improve. Notably, even with the latest DeepSeek-R1-Distill-Qwen-7B, a state-of-the-art reasoning model with built-in self-evaluation and self-correction, ID-sampling consistently outperforms vanilla sampling. This is because while stronger models excel at self-correction, they remain suboptimal at determining when to initiate the self-correction process. Compared to earlier models, each trigger sentence has a more pronounced effect, allowing ID-sampling to correct errors that might otherwise persist without explicit intervention.
5.2.2 AIME-24
Phi-4
@@ 5.2.2 AIME-24
Llama
$N$ Vanilla ID-Sampling
BoN cons@N BoN cons@N
1 0.00 0.00 3.45 3.45
4 0.00 0.00 0.00 3.45
8 0.00 0.00 0.00 3.45
16 3.45 0.00 0.00 3.45
32 3.45 0.00 10.34 6.90
Table 1: Results for different models with different number of samples on AIME-24 dataset. Top: Phi-4 comparison. Bottom: Llama-3.1-1B-Instruct comparison (BoN vs. ID-BoN). Best results in each setting is highlighted in bold. Here, the number of samples $N$ is not calculated as equivalent $N$.
Figure 5: BoN results for DeepSeek-R1-Distill-Qwen-7B on AIME with different $\gamma$, compared to the vanilla sampling. The numbers in the bracket represent the $\gamma$ used in ID-sampling. Given the diverse time used for different $\gamma$, the x-axis is the actual $N$ used rather than the equivalent one.
Figure 6: cons@N results for DeepSeek-R1-Distill-Qwen-7B on AIME with different $\gamma$, compared to the vanilla sampling. The numbers in the bracket represent the $\gamma$ used in ID-sampling. Given the diverse time used for different $\gamma$, the x-axis is the actual $N$ used rather than the equivalent one.
Method Relative Time
Vanilla 1.00
ID(2.5) 1.17
ID(2.0) 1.88
ID(1.5) 2.29
ID(1.2) 3.53
Table 2: The relative time cost for ID-sampling for different $\gamma$ from Deepseek-R1-distill-Qwen-7B on AIME-24.
The AIME-24 dataset is a recently introduced, highly challenging benchmark that is explicitly excluded from model training data. In this setting, reward models exhibit significantly lower accuracy, and the pass rates of various LLMs are notably lower compared to MATH-500 (Zhang et al., 2025). This makes AIME-24 a more rigorous test of a model’s reasoning and generalization capabilities.
We present our results in Figures 5, 6, and Table 1. Notably, for Best-of-N (BoN) sampling, we observe a counterintuitive trend: performance decreases as $N$ increases and more samples are collected. This decline is attributed to false comparisons introduced by the reward model, which may misrank responses. However, even at the largest $N$ setting (i.e., $N=32$), BoN still underperforms compared to ID-sampling across all three models.
This result underscores the effectiveness of ID-sampling on challenging datasets like AIME, where the self-correction mechanism significantly enhances the model’s reasoning and iterative refinement capabilities. On the other hand, due to the intrinsic difficulty of the dataset, we do not observe a substantial benefit from scaling $N$ in the current setting. Specifically, if an LLM can solve a given problem, it tends to succeed consistently; conversely, if it fails, generating more responses does not improve accuracy. This finding highlights the necessity of generating more diverse responses, even when self-correction mechanisms are incorporated.
@@
Given the significant performance gains of ID-sampling on AIME, we conducted an ablation study to analyze the impact of the scaling factor $\gamma$ on ID-sampling. We present the pass rate results in Figures 5 and 6, and report the relative inference time for each setting in Table 2.
Our find that adjusting $\gamma$ substantially affects both performance and computational cost. The worst-performing setting was $\gamma=2.5$, which failed to provide a stable and significant improvement over vanilla sampling. The best performance was achieved at $\gamma=1.2$, where self-evaluation and self-correction were triggered frequently. We also tested smaller values $\gamma=0.8,1.0)$, but these settings were computationally infeasible, failing to complete within 10× the runtime of vanilla sampling, and were thus omitted.
Interestingly, we observed a non-convex relationship between $\gamma$ and performance, with the second-best setting being $\gamma=2.0$. This non-convex behavior complicates hyperparameter selection. From our perspective, although increasing the number of responses does not provide a clear scaling advantage in the current setting, runtime efficiency remains a critical factor. Notably, selecting $\gamma=1.2$ nearly doubles the runtime compared to $\gamma=2.0$. Given this trade-off, we adopt $\gamma=2.0$ as the default setting, as it balances inference speed and output quality.
@@ 6 Discussions
Due to computational constraints and the sudden release of the DeepSeek-R1 series, which introduced new phenomena that deserve investigation, our study primarily focuses on ID-sampling without integrating it into tree-search algorithms. However, we emphasize that our proposed framework can be directly extended to methods such as beam search and Monte Carlo Tree Search (MCTS). In these cases, the budget in Algorithm 1 can be directly interpreted as the computational budget in MCTS or the number of iterations in beam search. One challenge in this extension is that self-correction mechanisms reduce the reliance on early model outputs being correct, distinguishing the need for a new generation of PRMs from previously released ones. However, future researchers who develop or have access to more accurate PRMs that explicitly model the self-correction process can leverage our ID-sampling framework to enhance the reasoning capabilities of fixed models.
Beyond combining ID-sampling with tree-search algorithms, several promising research directions could further enhance performance. However, these avenues remain largely unexplored due to the substantial computational resources required. First of all, in our study, we employ a fixed trigger sentence—“Wait! Maybe I made some mistakes! I need to rethink from scratch.”—as introduced by Zhao et al. (2024). However, recent findings from DeepSeek-R1 suggest that self-evaluation and self-correction can emerge spontaneously without explicit manual prompting, implying that trigger sentences can be highly flexible. Moreover, DeepSeek-R1 models demonstrate a tendency to self-correct locally, often requiring only minimal prompts—such as a single word (“Wait!”)—followed by context-specific revisions. This suggests a significant opportunity: modifying the trigger sentence could substantially impact both the computational efficiency and the pass rate performance of ID-sampling. In our preliminary experiments, we tested a minimal trigger sentence—“Wait!”—but found that it was insufficient to consistently induce self-correction, often introducing noise rather than meaningful refinements. As reasoning models like o1 continue to evolve, future research could provide deeper insights into optimizing trigger sentences—both in form and function—to maximize the effectiveness of self-reflection mechanisms. A second promising research direction is exploring alternative frequencies for inserting trigger sentences beyond the current geometric sequence approach. While ID-sampling provides a key advantage—offering a theoretical guarantee of bounded additional computational cost compared to one-shot generation—it remains possible that alternative designs, even without strong theoretical guarantees, could yield superior empirical performance. However, the design space for such strategies is vast, making exhaustive exploration infeasible. We believe that our proposed algorithm establishes a strong baseline for future research aimed at optimizing the frequency of self-evaluation and self-correction triggers, paving the way for more efficient and effective reasoning mechanisms. Lastly, in the early stages of our project, we investigated whether the initial budget impacts the performance of ID-sampling. However, we did not observe any consistent effect, particularly because the geometrically increasing budget quickly equalizes differences across initial budget settings after a few trigger sentences are introduced. Nevertheless, it remains possible that different models exhibit varying sensitivity to this hyperparameter, suggesting an avenue for further exploration.
@@
Our proposed method also has several clear limitations. The most significant is that, as highlighted in the latest DeepSeek-R1 technical report, self-correction capabilities can emerge spontaneously through reinforcement learning. This suggests that our algorithm has a more constrained range of applications. In particular, as discussed in the experimental section, the performance of our method varies significantly depending on the underlying model, and the improvement margin of ID-sampling is notably smaller on newer models. However, we argue that given the demonstrated effectiveness of using supervised fine-tuning (SFT) for distillation from existing models (DeepSeek-AI et al., 2025), along with theoretical guarantees that bound the additional computational cost, ID-sampling remains a valuable approach for efficiently generating data in certain scenarios. Moreover, incorporating trigger sentences into the generation process requires ID-sampling to invoke multiple generation steps. While this theoretically incurs no additional cost on the KV-cache, in practice, it can lead to increased inference time unless one modifies the inference engine—such as vLLM—to manually store and reuse the KV-cache. Additionally, due to the need for multiple sampling steps, ID-sampling introduces extra assistant tokens during generation, which may cause a slight distribution shift when applied to black-box models. However, our experiments indicate that the impact of this shift is generally minimal and can be ignored in most cases.

### v2

@@ 4.3 Theoretical Analysis
A key challenge in ID-sampling is that budget control occurs before each manually triggered self-evaluation and self-correction step without explicitly analyzing the actual generated responses. This can lead to unnecessary iterations, potentially increasing computational costs. However, due to the design of our ID-sampling method, we establish important theoretical guarantees. In particular, we provide a bound on the total number of tokens generated before reaching the final answer, as formalized in the following theorem.
@@
Suppose the final answer obtained through ID-Sampling needs a budget of $L$ in normal sampling without manual injection. Then the total number of budget used is no more than $\frac{\gamma*L}{\gamma-1}$.
@@
Note that the budget for each generation iteration follows a geometric sequence with common ratio $\gamma$. The theorem follows directly from a summation of this geometric series.
The theorem does not guarantee the quality of the generated answer or the generation distribution to be the same; rather, it ensures that our method introduces no significant additional computational overhead during the generation of a single response. And we will later show that this overhead will be even smaller in practice. The observed superiority of our algorithm stems from the intuition that explicitly injecting a trigger sentence biases the response distribution toward higher-quality outputs.
@@ 5.1 Experiment Setup
In our experiments, we evaluate ID-sampling on a set of mathematical problems. Given the strong problem-solving capabilities of modern models, which inherently possess self-evaluation and self-correction mechanisms, we focus on more challenging benchmarks and omit relatively easier datasets like GSM8K (Cobbe et al., 2021). Instead, we directly assess performance on competition-level and Olympiad-level benchmarks, which are more commonly included in technical reports of recently released models, including MATH-500 (Lightman et al., 2023), AIME-24 and AIME-25 (of America, 2025). Specifically, MATH-500 is a curated subset of 500 problems from the original MATH dataset. Meanwhile, AIME (American Invitational Mathematics Examination) is a highly competitive exam aimed at identifying top-performing high school students in the U.S. The AIME-24 and AIME-25 dataset each comprises 30 problems sourced from the AIME I and II exams, providing a rigorous benchmark for assessing reasoning capabilities under competition-level constraints. For both datasets, answers are evaluated using symbolic-based checkers to ensure that all mathematically equivalent responses are recognized as correct.
For our experiments, we use Llama-3.1-1B-Instruct (Dubey et al., 2024) and Phi-4 (Abdin et al., 2024) as examples for non-reasoning models, and DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-32B(DeepSeek-AI et al., 2025) and Qwen3-8B for reasoning models. For reward models for BoN, we employ Qwen-2.5-Math-PRM-7B (Zhang et al., 2025), which demonstrates superior accuracy compared to other available models. This choice ensures a robust evaluation of reasoning performance in our study.
For baselines, we compare our approach against vanilla sampling without any manual injection. For evaluation, we report three key pass rate metrics across the datasets:
@@
    Pass@N - The pass rate, which measures whether or not at least one of the N total responses is correct. We use an unbiased estimation, following the calculation in (Chen et al., 2021).
  3. 3.
@@
For both BoN and majority voting, a single aggregated answer is compared against the ground-truth solution to measure accuracy. We report BoN on Math500 datasets and for non-reasoning models on AIME datasets, given that its response length is within the context length limit of the reward models, and report Pass@n for reasoning models on AIME datasets because the normal response length is ¿12K and beyond the context length of popular reward models.  [^1] We have tried to use techniques like Yarn to extend the context length limit to a longer context. However, the model will lead to a consistent decrease rather than an increase when we increase the number of responses, and thus, we decided to just report Pass@n on longer context answers.
For trigger sentence, we choose to use a whole sentence ”Wait! Maybe I made some mistakes! I need to rethink from scratch.” for non-reasoning models and a single word ”wait” for reasoning models. We will provide a discussion about this choice later. By default, we use $\gamma=2.0$ for ID-sampling, and we will later provide ablation study results on this. Due to the page limit, we leave other hyperparameters in the appendix.
@@ 5.2.1 Math-500
(a) Llama-3.1-1B-Instruct
(b) Phi-4
(c) Deepseek-R1-Distill-Qwen-7B
Figure 3: Math-500 dataset: Pass rate results for different models. The x-axis is the equivalent N after considering the extra time used by ID-sampling.
Before evaluating the pass rate, we first analyze the runtime overhead of ID-sampling on Math-500 datasets. We observe that ID-sampling incurs approximately 1.6–1.9× the total wall-clock time compared to the baseline of vanilla sampling for non-reasoning models, and 1.1–1.3× for reasoning models. Given that this additional computational cost could instead be allocated to generating more responses and selecting the best one, we present our results in terms of an equivalent N. Specifically, if the original results correspond to Bo8, we report them as equivalent $N=16$, as ID-sampling consistently completes within twice the runtime of the original method.
The MATH-500 dataset is a curated subset of the MATH dataset. Originally used by OpenAI to train process reward models (PRMs) (Lightman et al., 2023), this dataset benefits from the high accuracy of recently released reward models like Qwen-2.5-Math-PRM (Zhang et al., 2025). As a result, in this setting, Best-of-N (BoN) sampling can serve as a reliable sampling method even in the absence of a ground-truth checker. We present our results for different models on Math-500 in Fig. 3(a), 3(b), 3(c). We observe a notable difference in the efficacy of ID-sampling for non-reasoning models based on the aggregation method. While ID-sampling yields performance gains with increasing sample size (N) under a Best-of-N (BoN) selection strategy, it consistently performs poorly when results are combined via majority voting. In the meantime, ID-sampling stably outperforms regular sampling even when equivalent N is considered. Our results also indicate that the performance gap between ID-sampling and vanilla sampling increases as model capabilities improve. Notably, even with the DeepSeek-R1-Distill-Qwen-7B, a strong reasoning model with built-in self-evaluation and self-correction, ID-sampling consistently outperforms vanilla sampling. This is because while stronger models excel at self-correction, they remain suboptimal at determining when to initiate the self-correction process. Compared to earlier models, each trigger sentence has a more pronounced effect, allowing ID-sampling to correct errors that might otherwise persist without explicit intervention.
5.2.2 AIME
$N$ Vanilla ID-sampling
BoN cons@N BoN cons@N
1 0.00 0.00 3.45 3.45
4 0.00 0.00 0.00 3.45
8 0.00 0.00 0.00 3.45
16 3.45 0.00 0.00 3.45
32 3.45 0.00 10.34 6.90
(a) Llama-3.1-1B-Instruct
@@ 5.2.2 AIME
(b) Phi-4
Table 1: AIME-24 dataset: Pass rate results for different models with different number of samples. The best results in each setting are highlighted in bold. Since performance saturates quickly on these weaker LLM models, meaning that even the highest pass rate with $N=32$ does not exceed the $N=1$ pass rate for ID-sampling, we omit the use of equivalent N in this setting.
Figure 4: Qwen3-8B results on AIME-24.
Figure 5: Qwen3-8B results on AIME-25.
Figure 6: Deepseek-R1-distill-Qwen-7B results on AIME-25.
Pass cons
Vanilla 72.6 83.3
ID-sampling 75.5 86.7
Table 2: Pass@1 and cons@32 comparison between Vanilla sampling and ID-sampling on Deepseek-R1-distill-Qwen-32B on AIME-24 dataset.
Vanilla ID(2.5) ID(2.0) ID(1.5)
Rel. Time 1.00 1.07 1.09 1.39
Table 3: Relative wall-clock time cost of ID-sampling for varying $\gamma$ values using DeepSeek-R1-Distill-Qwen-7B on the AIME-24 dataset.
(a) Pass@N
(b) cons@N
Figure 7: AIME-24 Dataset: Pass rate results for DeepSeek-R1-Distill-Qwen-7B with different $\gamma$ for ID-sampling, compared to vanilla sampling. The numbers in brackets are the $\gamma$ used in ID-sampling. As different values of $\gamma$ result in significantly varying runtimes, as shown in Table 3, we again omit the use of equivalent $N$ in this analysis.
The AIME datasets, including both AIME-24 and AIME-25, are highly challenging benchmarks that are now widely used as challenging datasets for reasoning models (DeepSeek-AI et al., 2025; Huang et al., 2024). In this setting, reward models exhibit significantly lower accuracy, and the reasoning is generally longer.
We first present the results for non-reasoning models in Table 1. We observe that ID-sampling can help improve the pass rate, and because of the fast saturation of pass rate, ID-sampling can even help the model to surpass its previous ceiling, i.e., $N=1$ with ID-sampling is better than $N=32$ with vanilla sampling.
For reasoning models, we first observe that the runtime difference between vanilla sampling and ID-sampling becomes negligible. For Qwen3 models, the time ratio remains within $1\pm 0.02$, and for R1-distill models, it is within $1\pm 0.1$ with $\gamma=2.0$. As a result, the use of equivalent $N$ has little practical effect in this context. We present results on reasoning models in Fig. 4, 5, 6, 7, and Table 2. Across different models on the two latest AIME datasets, ID-sampling consistently outperforms vanilla sampling. Specifically, on both datasets, Qwen3-8B achieves a larger improvement margin compared to Deepseek-R1-distill-Qwen-7B. These results highlight not only the robustness of our algorithm on recent models and benchmarks, but also the promise of ID-sampling as models gain stronger built-in self-correction capabilities and stronger reasoning capabilities.
@@
Given the stable performance gains of ID-sampling on AIME, here we conducted an ablation study to analyze the impact of the scaling factor $\gamma$ on ID-sampling. We present the pass rate results for DeepSeek-R1-Distill-Qwen-7B on AIME-24 in Figures 7(a) and 7(b), and report the relative inference time for each setting in Table 3.
We find that adjusting $\gamma$ significantly impacts both performance and computational cost. In terms of runtime, $\gamma=1.5$ yields the highest cost, whereas the other two settings remain within $\times1$ the wall-clock time of vanilla sampling. Regarding performance, $\gamma=2.0$ consistently achieves the best results in terms of Pass@$N$, while $\gamma=2.5$ occasionally outperforms in terms of Cons@$N$, though with a small margin. Overall, ID-sampling proves to be a more effective sampling strategy than vanilla sampling, provided that $\gamma$ is not too small, which would undermine our goal of emphasizing early-stage trigger sentence injection.
Interestingly, we observed a non-convex relationship between $\gamma$ and performance even without considering the time used, with the second-best setting being $\gamma=2.0$. This non-convex behavior complicates hyperparameter selection. However, as we described in our motivation section, we want our algorithm to inject more in the early stage and drop fast in later stage, and thus recommend using $\gamma=2.0$ given its balance between performance and time.
@@ 6 Discussions
Choice of Trigger Sentence
In our experiments, we use different trigger sentences for non-reasoning models and reasoning models. The difference is caused by the nature of the models. For non-reasoning models, a single word ”wait” cannot trigger the self-correction process and can only introduce noise to the generation. For reasoning models, we found that for unknown reason, adding the whole sentence will trigger the generation of an end-of-think token for a certain probability, and leads to early stop in the generation sequence. While this does not always harm the performance, this is not something we want, as we want to use the trigger sentence to trigger self-correction. Overall, we believe the choice of trigger sentence should be case-by-case, depending on the underlying model and the specific usage of the algorithm. But in principle, a single word, ”wait,” will be sufficient for most reasoning models.
Limitations
@@
Our proposed method also has several clear limitations. The most significant is that the performance of our method could significantly depend on the underlying model. While we have used our experiments to show that ID-sampling works on popular models, it is possible for some models to be incompatible with the algorithm, especially if the models are trained with losses that incorporate more than just the accuracy of the final answer. Moreover, incorporating trigger sentences into the generation process requires ID-sampling to invoke multiple generation steps. While this theoretically incurs no additional cost on the KV-cache, in practice, it can lead to increased inference time unless one modifies the inference engines like vLLM and SGLang to manually store and reuse the KV-cache. To address this concern, we use equivalent $N$ in our experiments, which accounts for total wall-clock time, to enable a fair comparison across different sampling methods. Additionally, due to the need for multiple sampling steps, ID-sampling is currently tested only on open-source models. While the same idea can also be applied to black-box models through multi-round generations, this will introduce extra assistant tokens during generation, which may cause a slight distribution shift when applied to black-box models. However, our experiments indicate that the impact of this shift is generally minimal and can be ignored in most cases.